Unit conversion for Soluble Reactive Phosphorus (SRP PO4-P)

Convert the following:

  1. 9706.7 (mg/m2) to (μg 1-1)
  2. 6.4 (μg 1-1) to (mg/m2)
  3. 0.04 mg/l to (μg 1-1)
1
Expert's answer
2022-05-20T13:25:05-0400

mg/m2 is weight by area that is pressure. But mg/l is weight by volume that is density. Pressure and density are not convertible.

The units you wish to convert are NOT compatible. You have to use “mg in CUBIC meters”, which you can convert to mg in Liters.

I think there is a typo here, then

1. 9706.7 (mg/m3) = 9706.7 (μg/l)

2. 6.4 (μg/l) = 6.4 (mg/m3)

3. 0.04 (mg/l) = 40 (μg/l)